Changelog 19012010

BOM changes:
  • replaced IC1 (was: MC100EPT23DTG, TSSOP) with MC100EPT23DG (SO-8 package) due to problems with availability.
  • replaced IC3 (was: MC100LVEL29DG, SOIC16) with MC100EP29DTG (TSSOP) due to problems with availability, PCB space constraints and lack of VBB bias output.
  • replaced output buffer (IC11) with SN64BCT25244DW (the same as in CTRV/CTRP - single channel can drive 50 ohm load)
  • replaced IC18 (MC100EPT21 - calibration buffer) with MC100EPT23DG to reduce the number of different components in the design
  • replaced serial termination resistors next to the TDC with 8x resistor ladders due to space reasons
Schematic changes:
  • delay chain flip flops now use VBB to bias data inputs
  • routed individual LVPECL clocks to each output stage flip flop (with length matching) to reduce the load and avoid termination/biasing problems
  • LVDS clock to the FPGA (CLK_M2C) is now AC-coupled, as the PLL Vdd is 3.3 V and the IOB Vdd on the carrier is 1.8 V
  • added 100 nF decoupling capacitor for each input stage IC
  • PLL reset re-wired to RESETn (it's active 0, CERN libary components don't have dashes over inverted pins with default Altium settings)
  • reordered outputs in AD9516 PLL to improve routing
PCB changes:
  • general cleanup
  • transition from 4 layers -> 6 layers (was necessary for diff pair routing with length matching and to avoid any signals crossing power planes)
  • added guard ring (BYPASS) around the PLL filter
  • moved the voltage reference closer to the PLL
  • added thermal via patterns under AD9516 (dissipates a lot of heat)
  • added OHWR/CERN notice on the silkscreen
  • fixed mounting holes connections
  • updated stackup information (standard 1.55mm thickness, 18/35/35/35/35/18 um copper, 0.2/0.36/0.36/0.36/0.2 mm buildup).